OpenStudy (anonymous):

Solve using the elimination method. If the system has an infinite number of solutions, use set builder notation to write the solution set. If the system has no solution, state this. x + 9y = 79 -7x + 9y = 23

OpenStudy (anonymous):

eliminate y's by multiplying first equation by -1

OpenStudy (anonymous):

-x - 9y = -79 would that be right for the first one?

OpenStudy (anonymous):

GREAT!

OpenStudy (anonymous):

so you should get -8x=-56 right?

OpenStudy (anonymous):

how do you get that?

OpenStudy (anonymous):

so x=7 plug x into second eqn to get y

OpenStudy (anonymous):

the ys cancel and you have -x + -8x and -79 + 23

OpenStudy (anonymous):

sorry -x + -7x

OpenStudy (anonymous):

oh okay. so you use it to both equations after you solve the first one?

OpenStudy (anonymous):

you can pick one equation to plug back into to get the other variable...preferably the one you didn't alter

OpenStudy (anonymous):

plug both x and y into both original equations to check

OpenStudy (anonymous):

-7-9(8)=-79 -7-72=-79 -79=-79 TRUE

OpenStudy (anonymous):

OK. So what do I do after I plug it into the second equation?

OpenStudy (anonymous):

-7(7)+9(8)=23 -49+72=23 23=23 TRUE

OpenStudy (anonymous):

solve for y then combine (x,y) for solution (7,8)

OpenStudy (anonymous):

So, (7,8) would be the answer correct? That's what I got.

OpenStudy (anonymous):

GREAT!!!!!

OpenStudy (anonymous):

Thank you so much. How would I know if an equation has no solution or infinite solutions?

OpenStudy (anonymous):

no solution would give you different numbers on each side of equal side such as -2=0 and infinite solutions would give you same number on each side like 5=5

OpenStudy (anonymous):

parallel lines have no solution and same line have infinite solutions so if you put them in y=mx+b form and look at slope (m)-same for parallel

OpenStudy (anonymous):

Ok. thanks

OpenStudy (anonymous):

welcome...medal or fan me please :)